OpenStudy (anonymous):

Change the following problem in standard form using integers. Show work! y=4x-11

OpenStudy (anonymous):

The question is too vague. Specify.

OpenStudy (anonymous):

That's all it says I don't know how to specify it any more

OpenStudy (anonymous):

No worries. I get what they are asking for.

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (anonymous):

so here is the solution: so the goal is to reach the ax+by=c form. first you move the term 4x to the other side so it becomes -4x+y=-11

OpenStudy (anonymous):

thats basically it but if you want to simplify more as you should: you divide the whole equation by -1: so, it looks like this \[\frac{ -4x }{ -1 }+\frac{ y }{ -1 }=\frac{ -11 }{ -1 }\]

OpenStudy (anonymous):

Why is it -4?

OpenStudy (anonymous):

And -11?

OpenStudy (anonymous):

okay so you added -4x on both sides your original equation was y=4x-11 after you add the -4x, you get: -4x+y=4x-4x-11 which is : -4x+y=0-11 and so -4x+y=-11

OpenStudy (anonymous):

is it still iffy?

OpenStudy (anonymous):

No that makes sense thanks :)

OpenStudy (anonymous):

yayy! have fun :)

OpenStudy (anonymous):

Is this one right? The equation is y= -2x-3 and I ended up with 3= -2x-y

OpenStudy (anonymous):

its correct except you wanna right it as -2x-y=3 they way you wrote it is not incorrect, but this is just the syntax that you will use later so better get a hang of it now

OpenStudy (anonymous):

Oh okay thanks and one last one haha. The equation is y=5/2x-22 and I got 2.5x+y= 22